# A Survey on Training-free Alignment of Large Language Models

### v2

@@ A Survey on Training-free Alignment of Large Language Models
Birong Pan1, Yongqi Li1, Weiyu Zhang2, Wenpeng Lu2, Mayi Xu1, Shen Zhou1, Yuanyuan Zhu1, Ming Zhong1, Tieyun Qian1,3thanks:   Corresponding author. 1School of Computer Science, Wuhan University, Wuhan, China 2Faculty of Computer Science and Technology, Qilu University of Technology, Shandong, China 3Zhongguancun Academy, Beijing, China {panbirong,qty}@whu.edu.cn
@@ 1 Introduction
Calculation [ DeRa Liu et al. (2024c), Proxy Tuning (Liu et al., 2024a),
Linear Alignment (Gao et al., 2024b), MOD (Shi et al., 2024),
@@ 3.2 In-Decoding TF Alignment
Several methods modify the logits of tokens during generation to align outputs with desired behaviors. Liu et al. (2024c, a); Shi et al. (2024); Fang et al. (2024) combine the logits of the alignment model and a reference model to guide generation, achieving FT-like effects without parameter updates. $\delta$ -UNLEARNING (Huang et al., 2024b) learns an offset from logits comparisons between smaller models, applying it to larger black-box LLMs to adjust their predictive behavior. Additionally, CoCA (Gao et al., 2024a) calibrates the output distribution by amplifying the model’s response to safety prompts. The method calculates logits difference before and after safety prompt insertion, ensuring robust alignment. Recently, IVG Liu et al. (2024e) uses implicit and explicit value functions to guide language model decoding at token and chunk-level respectively, efficiently aligning LLMs purely at inference time. Contrastive decoding has also been adopted. For example, MCA (Fu et al., 2024b) constructs expert and adversarial prompts to promote or suppress aligned targets dynamically. Lately, Linear Alignment (Gao et al., 2024b) relies on a novel parameterization for policy optimization under divergence constraints and estimates the preference direction using self-contrastive decoding.
@@
For guidance-based approaches, Qian et al. (2024); Wang et al. (2024b); Huang et al. (2024c); Bhattacharjee et al. (2024); Cao et al. (2025); Li et al. (2025b) extract guidance vectors to adjust the activation states of the target model during inference, achieving alignment. To predict next-token rewards for efficient and effective autoregressive generation, GenARM Xu et al. (2025) and CARDS (Li et al., 2025a) leverage the reward model to guide frozen LLMs toward expected distribution. PAD (Chen et al., 2024b) and ARGS (Khanov et al., 2024) leverage token-level rewards to guide the decoding process, dynamically guiding the base model’s predictions. Nudging Fei et al. (2024) and WSD (Song et al., 2025c) combine a large base model with a much smaller aligned model. RDS Zeng et al. (2025) designs a root classifier based on the discriminative capacity of queries, and then reorders the token and prioritizes the benign token. To directly align model outputs with human preferences, OPAD Zhu et al. (2025) guides the generation of a final output that aligns with the target principle in a token-by-token manner.
@@ 4.2 Future Directions
To compare various alignment methods more thoroughly, unified computational overhead metrics to evaluate memory usage, CPU/GPU utilization, and time taken for alignment steps need to be established. Additionally, cost-effective algorithms for input detection, logits adjustment, and response modification to reduce latency should be designed. For instance, HarmAug (Lee et al., 2024) distills a large teacher safety guard model into a smaller one. Also, it is necessary to investigate on-the-fly optimization strategies Zhu et al. (2025) that dynamically adapt computational resources based on complexity of tasks or efficiency requirements of users.
@@
Interpretability can help address this problem (Wu et al., 2024). It locates and edits fine-grained features, which can be used to align LLMs to human values. For example, by feeding prompts with negative and positive prefixes into LLMs, Leong et al. (2023) analyze internal contextualized representations to identify the toxicity direction of each attention head. However, many interpretability studies are primarily based on toy or theoretical models. Developing more practical interpretability TF alignment methods is necessary. Another way is to use principle-based guidance. OPAD Zhu et al. (2025) designs a principle-guided reward function to align model outputs with human preferences without FT. Moreover, considering the advantages of TF alignment in customization, explore the use of controllable TF alignment for personalized scenarios Zhu et al. (2025); Chen et al. (2024b).

### v4

@@ A Survey on Training-free Alignment of Large Language Models
Birong Pan1, Yongqi Li1, Weiyu Zhang2,thanks:   Corresponding authors., Wenpeng Lu2,∗, Mayi Xu1, Shen Zhou1, Yuanyuan Zhu1, Ming Zhong1, Tieyun Qian1,3,∗ 1School of Computer Science, Wuhan University, Wuhan, China 2Faculty of Computer Science and Technology, Qilu University of Technology, Shandong, China 3Zhongguancun Academy, Beijing, China {panbirong,qty}@whu.edu.cn
@@ 1 Introduction
Calculation [ DeRa Liu et al. (2024c), InRa Zhu et al. (2025b), MOD (Shi et al., 2024),
Linear Alignment (Gao et al., 2024b), Proxy Tuning (Liu et al., 2024a),
@@ 3.2 In-Decoding TF Alignment
Several methods modify the logits of tokens during generation to align outputs with desired behaviors. Liu et al. (2024c, a); Shi et al. (2024); Fang et al. (2024); Zhu et al. (2025b) combine the logits of the alignment model and a reference model to guide generation, achieving FT-like effects without parameter updates. $\delta$ -UNLEARNING (Huang et al., 2024b) learns an offset from logits comparisons between smaller models, applying it to larger black-box LLMs to adjust their predictive behavior. Additionally, CoCA (Gao et al., 2024a) calibrates the output distribution by amplifying the model’s response to safety prompts. The method calculates logits difference before and after safety prompt insertion, ensuring robust alignment. Recently, IVG Liu et al. (2024e) uses implicit and explicit value functions to guide language model decoding at token and chunk-level respectively, efficiently aligning LLMs purely at inference time. Contrastive decoding has also been adopted. For example, MCA (Fu et al., 2024b) constructs expert and adversarial prompts to promote or suppress aligned targets dynamically. Lately, Linear Alignment (Gao et al., 2024b) relies on a novel parameterization for policy optimization under divergence constraints and estimates the preference direction using self-contrastive decoding.
@@
For guidance-based approaches, Qian et al. (2024); Wang et al. (2024b); Huang et al. (2024c); Bhattacharjee et al. (2024); Cao et al. (2025); Li et al. (2025b) extract guidance vectors to adjust the activation states of the target model during inference, achieving alignment. To predict next-token rewards for efficient and effective autoregressive generation, GenARM Xu et al. (2025) and CARDS (Li et al., 2025a) leverage the reward model to guide frozen LLMs toward expected distribution. PAD (Chen et al., 2024b) and ARGS (Khanov et al., 2024) leverage token-level rewards to guide the decoding process, dynamically guiding the base model’s predictions. Nudging Fei et al. (2024) and WSD (Song et al., 2025c) combine a large base model with a much smaller aligned model. CAVGAN Li et al. (2025c) utilizes generative adversarial network to learn the security judgment boundary inside the LLM to achieve efficient alignment. RDS Zeng et al. (2025) designs a root classifier based on the discriminative capacity of queries, and then reorders the token and prioritizes the benign token. To directly align model outputs with human preferences, OPAD Zhu et al. (2025a) guides the generation of a final output that aligns with the target principle in a token-by-token manner.
@@ 4.2 Future Directions
To compare various alignment methods more thoroughly, unified computational overhead metrics to evaluate memory usage, CPU/GPU utilization, and time taken for alignment steps need to be established. Additionally, cost-effective algorithms for input detection, logits adjustment, and response modification to reduce latency should be designed. For instance, HarmAug (Lee et al., 2024) distills a large teacher safety guard model into a smaller one. Also, it is necessary to investigate on-the-fly optimization strategies Zhu et al. (2025a) that dynamically adapt computational resources based on complexity of tasks or efficiency requirements of users.
@@
Interpretability can help address this problem (Wu et al., 2024). It locates and edits fine-grained features, which can be used to align LLMs to human values. For example, by feeding prompts with negative and positive prefixes into LLMs, Leong et al. (2023) analyze internal contextualized representations to identify the toxicity direction of each attention head. However, many interpretability studies are primarily based on toy or theoretical models. Developing more practical interpretability TF alignment methods is necessary. Another way is to use principle-based guidance. OPAD Zhu et al. (2025a) designs a principle-guided reward function to align model outputs with human preferences without FT. Moreover, considering the advantages of TF alignment in customization, explore the use of controllable TF alignment for personalized scenarios Zhu et al. (2025a); Chen et al. (2024b).